OpenStudy (anonymous):

The arc corresponding to a central angle of 35 degrees in a circle of radius 10 feet measures _____ feet. Round your answer to two decimal places.

OpenStudy (whpalmer4):

Find the circumference of a circle with radius 10 feet. \(C = 2\pi r\) Find the fraction of that circumference that represents 35 degrees, if the entire circle is 360 degrees.

OpenStudy (anonymous):

\[C=20\pi\]

OpenStudy (whpalmer4):

Okay, now find the fraction of \(20\pi\) that represents 35/360 of the total.

OpenStudy (anonymous):

I don't know how to do that

OpenStudy (whpalmer4):

If I told you to find the fraction of \(20\pi\) that represents 1/2 of the total, could you do that?

OpenStudy (whpalmer4):

If you get the answer to this problem correct, you are paid $1000 by your teacher. My fee for helping you is 1/2 of what you get. How much will I get?

OpenStudy (anonymous):

$500

OpenStudy (whpalmer4):

Okay. If I get 3/4 of what you get, how much will I get?

OpenStudy (anonymous):

$750

OpenStudy (whpalmer4):

Right. So you understand the concept, whether you realize it or not. What is 35/360 of \(20\pi\)?

OpenStudy (anonymous):

6.19?

OpenStudy (whpalmer4):

Close: I get 6.1086. I wonder if you made a mistake while rounding? 6.1086 rounded to two decimal places would be 6.11. The 0 is the place where we are rounding, so we look one digit to the right. If that digit is 0-4, we keep the 0 we have in our rounding spot, otherwise we add 1 to the rounding spot's digit (and if that produces a carry, we have to propagate that to the left — 5.996 would round to 6.00 if rounding to two decimal places).

OpenStudy (anonymous):

Yea, I messed up the rounding.